next two terms of sequence 1,6,20,56
Answer by Edwin McCravy(20056)

       1,6,20,56

    Start with 1

               1
  1x2 +  4 =   6 
  6x2 +  8 =  20
 20x2 + 16 =  56
 56x2 + 32 = 144
144x2 + 64 = 352 

Edwin